A two-year financial instrument produces $129 in year t = 1 and $85 in year t = 2. Find the weight corresponding to t = 1, w1, in the definition of duration as a weighted average of the times at which the cash flows occur: D = 1*w1 + 2*w2. Assume the discount rate is 5% p.a
Answer #1

PV of CF1 = 128/1.05^1=121.9048

PV of CF2= 85/1.05^2 = 77.0975

Total = 199

W1= 121.9048/199 = 0.61258

Hence W1=0.61258